# Clarification and Coinduction of Tambara Functors

### v1

@@ A Tambara functor is a product-preserving functor from
$R:\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})\rightarrow\mathsf{Set}$
for which the commutative monoid constituting the value at any $G$-set is actually an abelian group.
@@
    $\mathrm{Res}_{L}^{H}\mathrm{Tr}_{K}^{H}=\prod_{g\in L\backslash H/K}\mathrm{Nm}_{L\cap gKg^{-1}}^{L}c_{g}\mathrm{Res}_{gLg^{-1}\cap K}^{K}$
@@
    ${R(X)}$${R(A)}$${X\times_{Y}\Pi_{f}A}$${Y}$${\Pi_{f}A}$$\scriptstyle{\mathrm{Nm}}$$\scriptstyle{\mathrm{Tr}}$$\scriptstyle{\mathrm{Res}}$$\scriptstyle{\mathrm{Nm}}$$\scriptstyle{\mathrm{Tr}}$
@@
The first three axioms are encoded by the statement that norms, restrictions, and the multiplicative monoid structure on each $R(G/H)$ determines a semi-Mackey functor. Thus the only qualitatively new kind of assumption is the exponential formula, which will be studied in much greater detail in Section 3.1.
@@
Induction of an $H$-set to a $G$-set defines norm maps which make the Burnside Green functor into a Tambara functor. For $G=C_{p}$ the Burnside Tambara functor has the following Lewis diagram:
[…]
@@
$\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})(X,-)\mathrm{.}$
@@
$R\boxtimes\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})(,-)\mathrm{.}$
@@
There is a notion of bi-incomplete Tambara functor, in which we only ask for the existence of some norms and transfers, not necessarily all of them. Bi-incomplete Tambara functors were introduced by Blumberg and Hill in [BH22] (and their progenitor, incomplete Tambara functors, were introduced in [BH18], alongside the free polynomial algebras over incomplete Tambara functors whose definition is essentially the same as in Definition 2.34). For a compatible pair (a pair of subsets of the set of subgroups of $G$ which encode respectively the allowed transfers and norms), the associated category of bi-incomplete Tambara functors enjoys many of the same properties as the category of Tambara functors. Our strongest results only apply to Tambara functors as we require existence of all norms for many arguments. In some cases, our results also apply to bi-incomplete Tambara functors, especially those which are really only dependent on properties of the $G$-ring $R(G/e)$. However, for clarity, we will omit indicating which results are true for bi-incomplete Tambara functors.
2.4. Coinduction and other adjunctions
There are functors relating $G$-Mackey, $G$-Green, and $G$-Tambara functors for varying $G$, arising from functors relating categories of $G$-sets for varying $G$. Namely, if $H\subset G$, then the forgetful functor $\mathrm{res}_{H}^{G}$ from $H$-sets to $G$-sets has a left adjoint $\mathrm{ind}_{H}^{G}$. It sends an $H$-set $X$ to $G\times_{H}X$.
@@ 2.4. Coinduction and other adjunctions
$\mathrm{Res}_{H}^{G}T:=T\circ\mathrm{ind}_{H}^{G}\mathrm{.}$
@@
$\mathrm{Coind}_{H}^{G}T:=T\circ\mathrm{res}_{H}^{G}\mathrm{.}$
@@
Coinduction of Mackey functors is left adjoint to restriction.
@@
$\mathrm{Coind}_{H}^{G}:H{-}\mathsf{Mack}\rightarrow G{-}\mathsf{Mack}$
is lax monoidal and oplax monoidal.
@@
By, for example, [Cha24, Lemma 6.6], $\mathrm{Res}_{H}^{G}$ is strong monoidal. Now $\mathrm{Coind}_{H}^{G}$ is both left and right adjoint to $\mathrm{Res}_{H}^{G}$, hence is both lax monoidal and oplax monoidal. ∎
@@
$\mathrm{Coind}_{H}^{G}R:=\mathrm{Coind}_{H}^{G}\mathrm{FP}(R)(G/e)\cong\mathrm{Fun}(G/H,R)\mathrm{.}$
Note that $\mathrm{Fun}(G/H,R)$ is a ring via pointwise multiplication, and the $G$-action is given as follows. Choose canonical representatives $g\in G$ for each coset in $G/H$. Let $\gamma\in G$ and $g_{1}\in G$ be a canonical representative for the coset $g_{1}H$, and let $g_{2}$ be the canonical representative of $\gamma g_{1}H$, so $g_{2}^{-1}\gamma g_{1}\in H$. Then
@@
Definition 2.42.
If $H\subset G$ and $R$ is an $H$-Ring, Mackey, Green, or Tambara functor, then we define ${}^{g}H:=gHg^{-1}$ and ${}^{g}R$ the ${}^{g}H$-Ring, Mackey, Green, or Tambara functor given by restricting along the conjugation isomorphism ${}^{g}H\cong H$.
@@ 3. A product decomposition for Tambara functors
Most important $G$-Tambara functors are clarified (for a list of examples, see Proposition 4.9). On the other hand, coinductions are not clarified in general. In particular, by the main result of [SSW25], every Nullstellensatzian $G$-Tambara functor is coinduced, hence is not clarified. There are many examples of field-like $G$-Tambara functors which are not clarified by virtue of being coinduced [Wis24]. The full subcategory of $G$-Tambara functors spanned by the category of clarified $G$-Tambara functors should be thought of as a localization obtained by formally setting the images of all nontrivial coinduction functors equal to zero.
@@ 3.1. Consequences of the exponential formula
In this subsection we collect many important consequences of the exponential formula. The first of these lemmas seems to be well-known, but to the author’s knowledge the rest of the results of this subsection are new. In principle, most of our formulas should be recoverable from a very careful analysis of the formula described in [AKGH21, Theorem 8.2], in the sense that they are all very special cases of it.
The key idea underlining all the proofs of this subsection is due to Mazur [Maz13]. In a fixed-point Tambara functor, all exponential formulas are determined by the double coset formula. For example, the norm of a sum of elements in the $G/e$ level is given by
$\mathrm{Nm}_{e}^{H}(a+b)=\prod_{h\in H}\left(ha+hb\right)\mathrm{.}$
@@
Lemma 3.4 for example may be interpreted through this lens as the special case where $a$ and $b$ are orthogonal idempotents fixed by $H$, which implies $\mathrm{Nm}_{H}^{K}(a+b)=\mathrm{Nm}_{H}^{K}(a)+\mathrm{Nm}_{H}^{K}(b)$. Unfortunately, the best way to make this argument entirely rigorous amounts to a careful analysis of exponential diagrams of $G$-sets and the corresponding exponential axiom for Tambara functors, and the argument obfuscates the above intuition as the key idea.
@@
Consider the exponential diagram corresponding to the map $G/K\sqcup G/K\rightarrow G/K\rightarrow G/L$ which classifies $(x,y)\mapsto\mathrm{Nm}_{K}^{L}(x+y)$. The Tambara axioms imply that this norm of a sum is computed by first restricting along evaluation $G/K\sqcup G/K\leftarrow G/K\times_{G/L}\Pi_{f}(G/K\sqcup G/K)$ for $f$ the quotient $G/K\rightarrow G/L$, then norming along the projection onto the second factor, and finally transfering along the projection $\Pi_{f}(G/K\sqcup G/K)\rightarrow G/L$.
We begin by analyzing $\Pi_{f}(G/K\sqcup G/K)$. It is the set of pairs
$(gL,\sigma:\{g^{\prime}K|g^{\prime}L=gL\}\rightarrow G/K\sqcup G/K)$
for which $\sigma(g^{\prime}K)=g^{\prime}K$ holds in one of the two copies of $G/K$. In particular, $\sigma(g^{\prime}K)$ is completely determined by a choice of either left or right factor. We can therefore identify $\Pi_{f}(G/K\sqcup G/K)$ with the set of pairs
$(gL,\sigma:\{g^{\prime}K|g^{\prime}L=gL\}\rightarrow\{\mathrm{l},\mathrm{r}\})$
where “l” and “r” stand for “left” and “right” respectively. Via this identification, the $G$ action is given by
[…]
The stabilizer of $(gL,\sigma)$ must be contained in the stabilizer $gLg^{-1}$ of $gL\in G/L$. If $\sigma$ is constant, the stabilizer of $(gL,\sigma)$ is precisely $gLg^{-1}$. On the other hand, if the stabilizer of $(gL,\sigma)$ is equal to $gLg^{-1}$, then the stabilizer of $g^{-1}(gL,\sigma)$ is $L$. In this case, $\sigma$ is an $L$-equivariant map $L/K\rightarrow\{\mathrm{l},\mathrm{r}\}$ with respect to the trivial action on the codomain. Since $L$ acts transitively on the right-hand side, $\sigma$ must be constant. We have thus proven that the stabilizer of $(gL,\sigma)$ is a proper subconjugate of $L$ if and only if $\sigma$ is not constant.
Recall that we implicitly extend any Tambara functor to non-transitive $G$-sets in the union product-preserving way. Norms, transfers, and restrictions along arbitrary $G$-set morphisms are determined by this extension as the respective products of norms, transfers, and restrictions in the appropriate factors. Note that $G/K\times_{G/L}\Pi_{f}(G/K\sqcup G/K)$ contains two copies of $G/K$ given by the $G$-orbits of $(eK,(eL,\sigma))$ where $\sigma$ is constant. The composition of restriction along the map to $G/K\sqcup G/K$ followed by projection onto the two factors corresponding to these two copies of $G/K$ is the identity. The map $G/K\times_{G/L}\Pi_{f}(G/K\sqcup G/K)\rightarrow\Pi_{f}(G/K\sqcup G/K)$ is given on the two factors of $G/K$ corresponding to constant $\sigma$ by the quotient $G/K\sqcup G/K\rightarrow G/L\sqcup G/L$. The norm along this map therefore sends $(a,b)$ in the factors corresponding to $G/K\sqcup G/K$ to $(\mathrm{Nm}_{K}^{L}(a),\mathrm{Nm}_{K}^{L}(b))$.
The remaining $G$-orbits of $\Pi_{f}(G/K\sqcup G/K)$ are given by $G/L_{i}$, where $L_{i}$ is a proper subconjugate of $L$. Transfer along $\Pi_{f}(G/K\sqcup G/K)$ is given as the composition of a product of proper transfers (corresponding to quotients $G/L_{i}\rightarrow G/L$) followed by the sum over all factors (corresponding to a fold map $G/L^{\sqcup n}\rightarrow G/L$). We therefore obtain the formula
$\mathrm{Nm}_{K}^{L}(a+b)=\mathrm{Nm}_{K}^{L}(a)+\mathrm{Nm}_{K}^{L}(b)+\sum_{i}\mathrm{Tr}_{L_{i}}^{L}(x_{i})$
where $L_{i}$ is a proper subconjugate of $L$ and $x_{i}\in R(G/L_{i})$ (in fact, $x_{i}$ is a product of norms of conjugates of restrictions of $a$ and $b$, although we do not require this). Induction supplies the general result. ∎
@@
As in Lemma 3.3, the exponential formula gives an alternate description of the norm of a sum. As before, we have
$\Pi_{f}(G/e\sqcup G/e)\cong(gH,gH\rightarrow\{\mathrm{l},\mathrm{r}\})$
with the same description of the $G$-action and stabilizers as in Lemma 3.3.
Observe that $G/e\times_{G/H}\Pi_{f}(G/e\sqcup G/e)$ is the set of pairs $(g,gH\rightarrow\{\mathrm{l},\mathrm{r}\})$, hence is a disjoint union of copies of $G/e$, as $G$ acts diagonally. Each orbit has a unique representative given by the tuple with $e$ in the first coordinate. The restriction is a product of projections: it is a map
@@
The norm along the projection onto $\Pi_{f}(G/e\sqcup G/e)$ is the product of a collection of norm and multiplication maps. Two factors in this product description respectively send the elements $a$ and $b$ in the factors of $R(G/e)$ corresponding to the two constant functions $\sigma$ to $\mathrm{Nm}_{e}^{H}(a)$ and $\mathrm{Nm}_{e}^{H}(b)$, as in the proof of Lemma 3.3. To conclude the proof, it suffices to show that every other coordinate of the image of the norm along $G/e\times_{G/H}\Pi_{f}(G/e\sqcup G/e)\rightarrow\Pi_{f}(G/e\sqcup G/e)$ is zero.
@@
$(g,\sigma:gH\rightarrow\{\mathrm{l},\mathrm{r}\})\mapsto(gH,\sigma:gH\rightarrow\{\mathrm{l},\mathrm{r}\})\mathrm{.}$
@@
The stabilizer of $gd$ is a conjugate of $H$, and by assumption there is an element $\gamma\in L$ which is not contained in this conjugate. Then we have $\gamma gd\neq gd$, hence $(\gamma gd)(gd)=0$. The norm $\mathrm{Nm}_{e}^{L}$ factors through $L$-orbits, because if $\rho\in L$, then right multiplication by $\rho$ defines a $G$-equivariant morphism $G/e\rightarrow G/e$ over $G/L$ (and the $G$-action on $R(G/e)$ arises from the isomorphism $G\cong\mathrm{Aut}_{\mathrm{Span}(G-\mathsf{Set}^{\mathrm{fin}})}(G/e)$ given by right multiplication). Thus we compute
@@ 3.2. The product decomposition
The functor $R\mapsto R(G/e)$ from $G{-}\mathsf{Tamb}$ to $G{-}\mathsf{Ring}$ reflects products in the sense that if $R(G/e)\cong S_{1}\times S_{2}$ is an isomorphism of $G$-rings, then there is a canonical isomorphism $R\cong R_{1}\times R_{2}$ of Tambara functors with $R_{i}(G/e)=S_{i}$ given on the bottom level by the specified isomorphism $R(G/e)\cong R_{1}(G/e)\times R_{2}(G/e)$.
@@
Combining Proposition 3.7 and Proposition 3.6, we are therefore reduced to showing that if $k$ is a Tambara functor with $k(G/e)$ coinduced from an $H$-ring, then $k$ is coinduced as a Tambara functor. We start by showing that this is true after passing to restriction.
Lemma 3.8.
Let $G$ be a finite group. Let $k$ be a $G$-Tambara functor such that $k(G/e)\cong\mathrm{Coind}_{H}^{G}R$ for some $H$-ring $R$. Then there exists a map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ for some $H$-Tambara functor $\ell$ such that the induced map $\mathrm{Res}_{H}^{G}k\cong\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}\ell$ is an isomorphism.
The proofs of Lemma 3.8 and Proposition 3.9 are by induction on the order of $G$. In particular, in the proof of Lemma 3.8, we will need to assume that Proposition 3.9 holds with $G$ replaced by some of its proper subgroups. Since the proof of Proposition 3.9 does not require the statement of Lemma 3.8 to hold for any groups of order strictly larger than $|G|$, this does not present any circularity issues.
Additionally, the proof simplifies enormously if $H$ is normal in $G$. In this case, $\mathrm{Res}_{H}^{G}k$ carries a $G/H$ action by permuting cosets in $G/H$ according to right multiplication. The fixed points of this action are an $H$-Tambara functor $\ell$ which one may straightforwardly show satisfies the desired properties. In fact the proof of Lemma 3.8 is a careful unpacking of this idea, abstracting it enough that the requirement that $H$ is normal disappears.
@@
We induct on the order of $G$. If $H=G$, there is nothing to prove, as $\mathrm{Res}_{G}^{G}$ and $\mathrm{Coind}_{G}^{G}$ are the identity functors. Now assume $H$ is a proper subgroup of $G$. Lemma 2.43 supplies an $H$-ring isomorphism
$\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}R\cong\prod_{g\in H\backslash G/H}\mathrm{Coind}_{H\cap{}^{g}H}^{H}\mathrm{Res}_{H\cap{}^{g}H}^{{}^{g}H}{}^{g}R\mathrm{.}$
The formula $\left(\mathrm{Res}_{H}^{G}k\right)\left(H/e\right)\cong\mathrm{Res}_{H}^{G}\left(k(G/e)\right)$, Proposition 3.6, and our inductive usage of Proposition 3.9 imply that we have an isomorphism
@@
where the middle map is the identity double coset factor of the restriction of the map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ and the last map is induced by conjugation by $g^{-1}$ in $\mathrm{Coind}_{H}^{G}\ell$. In fact, each of these maps is invertible: the middle map is clearly invertible, and the first and third maps have inverse induced by conjugation by $g^{-1}$ and $g$ respectively in $k$ and $\mathrm{Coind}_{H}^{G}\ell$. ∎
Proposition 3.9.
Proof.
Let $\ell$ be the $H$-Tambara functor of Lemma 3.8. Then we have a map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ which becomes an isomorphism after applying $\mathrm{Res}_{H}^{G}$. Since $\mathrm{Res}_{H}^{G}$ is computed by precomposing a Tambara functor with $\mathrm{Ind}_{H}^{G}$ and $\mathrm{Ind}_{H}^{G}H/L\cong G/L$, we see that $k(X)\rightarrow\mathrm{Coind}_{H}^{G}\ell(X)$ is an isomorphism whenever $X$ is a $G$-set for which all elements of $X$ have isotropy sub-conjugate to $H$.
@@
This implies a strengthening of one of the main theorems of [Wis24]. Namely, the corollary below was proven only in the case $G=C_{p^{n}}$. It reduces the study of all field-like Tambara functors to those whose bottom level is a field, since this is equivalent to being clarified for field-like Tambara functors.
@@
If $k$ is any field-like Tambara functor, then $k(G/e)\cong\mathrm{Coind}_{H}^{G}\mathbb{F}$ for some field $\mathbb{F}$ with an action of $H$ (as observed by David Chan and Ben Spitz, this is a straightforward consequence of Nakaoka’s criteria [Nak12]). Therefore $k\cong\mathrm{Coind}_{H}^{G}\ell$ for some $H$-Tambara functor $\ell$ with $\ell(H/e)\cong\mathbb{F}$. Since the restrictions in $k$ are all injective, the restrictions in $\ell$ are all injective. Thus by Nakaoka’s criteria, $\ell$ is field-like. It is visibly clarified. ∎
@@
Let $R$ be a $G$-Tambara functor such that either
  1. (1)
    the ring $R(G/e)$ is Noetherian, or
  2. (2)
    $R$ is domain-like (see Definition 5.21)
Then
@@
If $R(G/e)$ is Noetherian, Proposition 3.7 implies that we have a $G$-ring isomorphism
@@
where $T_{H}$ is an $H$-Tambara functor such that $T_{H}(H/e)\cong\mathrm{Coind}_{H}^{G}S_{H}$. Finally, Proposition 3.9 implies
@@
Suppose $H$ and $H^{\prime}$ are conjugate. Then $\mathrm{Coind}_{H}^{G}R_{H}\cong\mathrm{Coind}_{H^{\prime}}^{G}\mathrm{Res}_{H^{\prime}}^{H}R_{H}$ where the restriction is along a $G$-set isomorphism $G/H\cong G/H^{\prime}$. Since $R_{H}$ is clarified if and only if $\mathrm{Res}_{H^{\prime}}^{H}R_{H}$ is, and coinduction commutes with products, we may rearrange and combine terms so that at most one representative from each conjugacy class of subgroups of $G$ appears in our product decomposition.
Finally, if $R$ is domain-like, then the result follows immediately from Proposition 5.27 (whose proof only requires Proposition 3.9). In particular, a domain-like Tambara functor may not be written nontrivially as a product of two Tambara functors, so only one nonzero term appears in the product decomposition. ∎
@@
In future work, the author will show that, for $k$-algebras $R$ and $S$, the $k$-algebra $R\times S$ is étale if and only if $R$ and $S$ are. By virtue of Theorem 3.11, if $k(G/e)$ is Noetherian, checking whether or not a finitely presented $k$-algebra is étale amounts to checking étaleness of a collection of morphisms into coinductions of clarified Tambara functors. For example, we obtain Theorem 4.15 below.

### v2

@@ A Tambara functor is a product-preserving functor from
$R:\mathrm{bispan}(G-\mathrm{Set}^{\mathrm{fin}})\rightarrow\mathrm{Set}$
such that the additive monoid $R(X)$ is actually a group for each finite $G$-set $X$.
@@
It will be necessary later to posses a detailed understanding of the dependent product. As a set, it consists of the set of pairs $(y,\sigma)$ where $y\in Y$ and $\sigma:f^{-1}(y)\rightarrow A$ such that the composition $f^{-1}(y)\xrightarrow{\sigma}A\rightarrow X$ is the inclusion $f^{-1}(y)\subset X$. The action of $G$ is defined by $g(y,\sigma)=(gy,g\sigma)$ where $(g\sigma)(x)=g\sigma(g^{-1}x)$. The map $\Pi_{f}A\rightarrow Y$ is given by $(y,\sigma)\mapsto y$.
@@
    $\mathrm{Res}_{L}^{H}\mathrm{Nm}_{K}^{H}=\prod_{g\in L\backslash H/K}\mathrm{Nm}_{L\cap gKg^{-1}}^{L}c_{g}\mathrm{Res}_{gLg^{-1}\cap K}^{K}$
@@
[…]
@@
Coinduction of an $H$-set to a $G$-set defines norm maps which make the Burnside Green functor into a Tambara functor. For $G=C_{p}$ the Burnside Tambara functor has the following Lewis diagram:
$\mathbb{Z}[x]/(x^{2}-px)$$\mathbb{Z}$$\scriptstyle{x\mapsto p}$$\scriptstyle{1\mapsto x}$$\scriptstyle{a\mapsto a+\frac{a^{p}-a}{p}x}$$\scriptstyle{C_{p}-\mathrm{trivial}}$
@@
$\mathrm{bispan}(G-\mathrm{Set}^{\mathrm{fin}})(X,-).$
@@
$R\boxtimes\mathrm{bispan}(G-\mathrm{Set}^{\mathrm{fin}})(,-).$
@@
There are well-known change-of-group functors relating $G$-Mackey, $G$-Green, and $G$-Tambara functors for varying $G$, arising from change-of-group functors relating categories of $G$-sets for varying $G$. Namely, if $H\subset G$, then the forgetful functor $\mathrm{res}_{H}^{G}$ from $H$-sets to $G$-sets has a left adjoint $\mathrm{ind}_{H}^{G}$. It sends an $H$-set $X$ to $G\times_{H}X$.
@@
$\mathrm{Res}_{H}^{G}T:=T\circ\mathrm{ind}_{H}^{G}.$
@@
$\mathrm{Coind}_{H}^{G}T:=T\circ\mathrm{res}_{H}^{G}.$
@@
Coinduction of Mackey functors is also left adjoint to restriction.
@@
$\mathrm{Coind}_{H}^{G}:H{-}\mathrm{Mack}\rightarrow G{-}\mathrm{Mack}$
is lax symmetric monoidal and oplax symmetric monoidal.
@@
By, for example, [Cha24, Lemma 6.6], $\mathrm{Res}_{H}^{G}$ is strong symmetric monoidal. Now $\mathrm{Coind}_{H}^{G}$ is both left and right adjoint to $\mathrm{Res}_{H}^{G}$, hence is both lax symmetric monoidal and oplax symmetric monoidal. ∎
@@
$\mathrm{Coind}_{H}^{G}R:=\mathrm{Coind}_{H}^{G}\mathrm{FP}(R)(G/e)\cong\mathrm{Fun}(G/H,R).$
Note that $\mathrm{Fun}(G/H,R)$ is a ring via pointwise multiplication, and the $G$-action is given as follows. Fix a choice of representatives $g\in G$ for each coset in $G/H$. Let $\gamma\in G$ and $g_{1}\in G$ be the chosen representative for the coset $g_{1}H$, and let $g_{2}$ be the chosen representative of $\gamma g_{1}H$, so $g_{2}^{-1}\gamma g_{1}\in H$. Then
@@
Notation 2.25.
If $H\subset G$ and $R$ is an $H$-Ring, Mackey, Green, or Tambara functor, then let $^{g}H:=gHg^{-1}$ and $^{g}R$ be the $^{g}H-$Ring, Mackey, Green, or Tambara functor given by restricting along the conjugation isomorphism $^{g}H\cong H$.
@@ 3. A product decomposition for Tambara functors
Most important $G$-Tambara functors are clarified (for a list of examples, see Proposition 4.9). On the other hand, coinductions are not clarified in general. In particular, by the main result of [SSW25], every Nullstellensatzian $G$-Tambara functor is coinduced, hence is not clarified. There are many examples of field-like $G$-Tambara functors which are not clarified by virtue of being coinduced [Wis25b]. The full subcategory of $G$-Tambara functors spanned by the category of clarified $G$-Tambara functors should be thought of as a localization obtained by formally setting the images of all nontrivial coinduction functors equal to zero.
@@ 3.1. Consequences of the exponential formula
In this subsection we collect many important consequences of the exponential formula. The key idea underlining all the proofs of this subsection is due to Mazur [Maz13]. In a fixed-point Tambara functor, all exponential formulas are determined by the double coset formula. For example, the norm of a sum of elements in the $G/e$ level is given by
$\mathrm{Nm}_{e}^{H}(a+b)=\prod_{h\in H}\left(ha+hb\right).$
@@
This may be seen directly from the formula of [AKGH21, Theorem 8.2]. ∎
@@
The exponential formula gives an alternate description of the norm of a sum. We have
$\Pi_{f}(G/e\sqcup G/e)\cong G/H\times\mathrm{Fun}(G/H,\{\mathrm{l},\mathrm{r}\})$
as $G$-sets.
Observe the isomorphism $G/e\times_{G/H}\Pi_{f}(G/e\sqcup G/e)\cong G/e\times\mathrm{Fun}(G/H,\{\mathrm{l},\mathrm{r}\})$, which is a disjoint union of copies of $G/e$, as $G$ acts diagonally. Each orbit has a unique representative given by the tuple with $e$ in the first coordinate. The restriction is a product of projections: it is a map
@@
The norm along the projection onto $\Pi_{f}(G/e\sqcup G/e)$ is the product of a collection of norm and multiplication maps. Two factors in this product description respectively send the elements $a$ and $b$ in the factors of $R(G/e)$ corresponding to the two constant functions $\sigma$ to $\mathrm{Nm}_{e}^{H}(a)$ and $\mathrm{Nm}_{e}^{H}(b)$. To conclude the proof, it suffices to show that every other coordinate of the image of the norm along $G/e\times_{G/H}\Pi_{f}(G/e\sqcup G/e)\rightarrow\Pi_{f}(G/e\sqcup G/e)$ is zero.
@@
$(g,\sigma:gH\rightarrow\{\mathrm{l},\mathrm{r}\})\mapsto(gH,\sigma:gH\rightarrow\{\mathrm{l},\mathrm{r}\}).$
@@
The stabilizer of $gd$ is a conjugate of $H$, and by assumption there is an element $\gamma\in L$ which is not contained in this conjugate. Then we have $\gamma gd\neq gd$, hence $(\gamma gd)(gd)=0$. The norm $\mathrm{Nm}_{e}^{L}$ factors through $L$-orbits, because if $\rho\in L$, then right multiplication by $\rho$ defines a $G$-equivariant morphism $G/e\rightarrow G/e$ over $G/L$ (and the $G$-action on $R(G/e)$ arises from the isomorphism $G\cong\mathrm{Aut}_{\mathrm{Span}(G-\mathrm{Set}^{\mathrm{fin}})}(G/e)$ given by right multiplication). Thus we compute
@@ 3.2. The product decomposition
The functor $R\mapsto R(G/e)$ from $G{-}\mathrm{Tamb}$ to $G{-}\mathrm{Ring}$ reflects products in the sense that if $R(G/e)\cong S_{1}\times S_{2}$ is an isomorphism of $G$-rings, then there is a canonical isomorphism $R\cong R_{1}\times R_{2}$ of Tambara functors with $R_{i}(G/e)=S_{i}$ given on the bottom level by the specified isomorphism $R(G/e)\cong R_{1}(G/e)\times R_{2}(G/e)$.
@@
Combining Proposition 3.7 and Proposition 3.6, we are therefore reduced to showing that if $k$ is a Tambara functor with $k(G/e)$ coinduced from an $H$-ring, then $k$ is coinduced as a Tambara functor.
Proposition 3.8.
The proof simplifies enormously if $H$ is normal in $G$. In this case, $\mathrm{Res}_{H}^{G}k$ carries a $G/H$ action by permuting cosets in $G/H$ according to right multiplication. One may straightforwardly show that the fixed points of this action give the desired $\ell$. The proof of Proposition 3.8 is a careful unpacking of this idea, abstracting it enough that the requirement that $H$ is normal disappears.
@@
We induct on the order of $G$. If $H=G$, there is nothing to prove, as $\mathrm{Res}_{G}^{G}$ and $\mathrm{Coind}_{G}^{G}$ are the identity functors. Now assume $H$ is a proper subgroup of $G$. The first half of our argument will consist entirely of constructing a map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ which becomes an isomorphism after applying $\mathrm{Res}_{H}^{G}$.
To start with, Lemma 2.26 supplies an $H$-ring isomorphism
$\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}R\cong\prod_{g\in H\backslash G/H}\mathrm{Coind}_{H\cap{}^{g}H}^{H}\mathrm{Res}_{H\cap{}^{g}H}^{{}^{g}H}{}^{g}R.$
The formula $\left(\mathrm{Res}_{H}^{G}k\right)\left(H/e\right)\cong\mathrm{Res}_{H}^{G}\left(k(G/e)\right)$, Proposition 3.6, and our inductive hypothesis imply that we have an isomorphism of $H$-Tambara functors
@@
where the middle map is the identity double coset factor of the restriction of the map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ and the last map is induced by conjugation by $g^{-1}$ in $\mathrm{Coind}_{H}^{G}\ell$. In fact, each of these maps is invertible: the middle map is clearly invertible, and the first and third maps have inverse induced by conjugation by $g^{-1}$ and $g$ respectively in $k$ and $\mathrm{Coind}_{H}^{G}\ell$.
At this point, we have constructed a map $k\rightarrow\mathrm{Coind}_{H}^{G}\ell$ which becomes an isomorphism after applying $\mathrm{Res}_{H}^{G}$. Since $\mathrm{Res}_{H}^{G}$ is computed by precomposing a Tambara functor with $\mathrm{Ind}_{H}^{G}$ and $\mathrm{Ind}_{H}^{G}H/L\cong G/L$, we see that $k(X)\rightarrow\mathrm{Coind}_{H}^{G}\ell(X)$ is an isomorphism whenever $X$ is a $G$-set for which all elements of $X$ have isotropy sub-conjugate to $H$.
@@
This implies a strengthening of one of the main theorems of [Wis25b]. Namely, Corollary 3.9 below was proven only in the case $G=C_{p^{n}}$. It reduces the study of all field-like Tambara functors to those for which every level is a field, since this is equivalent to being clarified for field-like Tambara functors.
@@
If $k$ is any field-like Tambara functor, then $k(G/e)\cong\mathrm{Coind}_{H}^{G}\mathbb{F}$ for some field $\mathbb{F}$ with an action of $H$ (as observed by David Chan and Ben Spitz, this is a straightforward consequence of Nakaoka’s criteria [Nak12] and the Chinese remainder theorem). Therefore $k\cong\mathrm{Coind}_{H}^{G}\ell$ for some $H$-Tambara functor $\ell$ with $\ell(H/e)\cong\mathbb{F}$. Since the restrictions in $k$ are all injective, the restrictions in $\ell$ are all injective. Thus by Nakaoka’s criteria, $\ell$ is field-like. It is visibly clarified. ∎
@@
Let $R$ be a $G$-Tambara functor such that either the ring $R(G/e)$ is Noetherian. Then
@@
Proposition 3.7 implies that we have a $G$-ring isomorphism
@@
where $T_{H}$ is a $G$-Tambara functor such that $T_{H}(G/e)\cong\mathrm{Coind}_{H}^{G}S_{H}$. Finally, Proposition 3.8 implies
@@
Suppose $H$ and $H^{\prime}$ are conjugate. Then $\mathrm{Coind}_{H}^{G}R_{H}\cong\mathrm{Coind}_{H^{\prime}}^{G}\mathrm{Res}_{H^{\prime}}^{H}R_{H}$ where the restriction is along a $G$-set isomorphism $G/H\cong G/H^{\prime}$. Since $R_{H}$ is clarified if and only if $\mathrm{Res}_{H^{\prime}}^{H}R_{H}$ is, and coinduction commutes with products, we may rearrange and combine terms so that at most one representative from each conjugacy class of subgroups of $G$ appears in our product decomposition. ∎